Consider a solenoid with an air core, with 3000 loops, with a measure of 70 cm, and with a diameter of 3 cm. If a 6 A current is sent through it, how much is the value of the magnetic formed within it?


1
Expert's answer
2022-03-20T18:51:24-0400

The magnetic field inside solenoid


"B=\mu_0 \frac{n}{\sqrt{d^2+l^2}}I""B=4\pi\times 10^{-7}\times \frac{3000/0.70}{\sqrt{0.03^2+0.70^2}}\times6=4.6\times 10^{-4}\:\rm T"